# Towards the Usage of Window Counting Constraints

### v2

@@ 2. Games with Counting Constraints
    For $p=EGO$, $CC_{min}(EGO,a,k,l)$ is defined as the abbreviation for “The player $EGO$ plays $a$ at least $k$ times out of $l$ of its own turns.”. $CC_{min}(EGO,a,k,l)$ is satisfied on $\pi$, if for all $i\in\mathbb{N}_{0}$ holds $|\{\sigma_{2m}\,|\,\sigma_{2m}\models a,\,i\leq m\leq i+l\}|\geq k$.
  3. (3)
    For $p=ALTER$, $CC_{max}(ALTER,a,k,l)$ is defined as the abbreviation for “The player $ALTER$ plays $a$ at most $k$ times out of $l$ of its own turns.”. $CC_{max}(ALTER,a,k,l)$ is satisfied on $\pi$, if for all $i\in\mathbb{N}_{0}$ holds $|\{\sigma_{2m+1}\,|\,\sigma_{2m+1}\models a,\,i\leq m\leq i+l\}|\leq k$.
  4. (4)
    For $p=ALTER$, $CC_{min}(ALTER,a,k,l)$ is defined as the abbreviation for “The player $ALTER$ plays $a$ at least $k$ times out of $l$ of its own turns.”. $CC_{min}(ALTER,a,k,l)$ is satisfied on $\pi$, if for all $i\in\mathbb{N}_{0}$ holds $|\{\sigma_{2m+1}\,|\,\sigma_{2m+1}\models a,\,i\leq m\leq i+l\}|\geq k$.
@@
  • “The player $EGO$ plays $y$ at least $1$ times out of $5$ of its own turns.” and
@@
The play $\pi=(1,\{y\},2,\{a\},3,\{x\},2,\{a\},3\{x,y\},4,\{a\})^{\omega}$ on the game graph satisfies all of the four window counting constraints, since each “sliding window” shows constraint satisfaction. Figure 2 illustrates the constraint fulfillment for the first two constraints.
The prefix $pre=(1,\{x\},4,\{a\},1,\{x\},4,\{x\},1,\{x\},4)$ of another play on the shown graph is not satisfying “The player $EGO$ plays $y$ at least $1$ times out of $5$ of its own turns.”, since no matter what $EGO$ does in its next move (even if it would be allowed to change the game graph for the next turns), it is not possible to extend prepre to a play that satisfies the constraint.
@@
  • Win is a winning condition for $EGO$ on the game graph $G^{\prime}$.
  • CCCC is a finite sets of counting constraints.
@@
Player $EGO$ wins a play on $G$ in which $ALTER$ satisfies its constraints, if Win is fulfilled in the play and the play satisfies all counting constraints on $EGO$. A strategy $\varphi$ of $EGO$ is winning for $EGO$ (or a winning strategy of $EGO$), if $EGO$ wins all plays in which $ALTER$ satisfies its constraints in $O(G,\varphi)$.
The fourth property of a game with counting constraints seems counter-intuitive, since it restricts the considered games to those in which $ALTER$ can basically ensure in each step to satisfy all of its constraints without any need of planning ahead. The main reason for introducing this property is as follows: We are not interested in plays that are only won by $EGO$ due to forcing $ALTER$ to loose due to violated counting constraints. Dismissing the property would lead to situations like depicted in Figure 3: The $EGO$-player does not have any counting constraints to obey but needs to avoid an unsafe state (colored in gray). $ALTER$ has the counting constraint ALTERALTER plays $b$ at least 1 time in 1 turn.” (hence, $ALTER$ needs to play $b$ in each of its turns). In the initial state, choosing to enter the upper sub-graph lets $EGO$ win. However, taking the transitions to the lower subgraph and consequently visiting an unsafe state would also lead to a win for $EGO$, since $ALTER$ is violating its constraint in the resulting play. This problem of winning by falsifying the assumptions, but not leading to intended behavior of the player is well-known [MPS19], [BEJK14].
@@
Figure 3. Graph for a safety game that cannot be used in a game with the counting constraint ALTERALTER plays $b$ at least 1 time in 1 turn.”. Circles represent states controlled by $EGO$, diamond-shaped states are controlled by $ALTER$. The state colored in gray marks an unsafe state. When the lower subgraph is entered, AA cannot fulfill its counting constraint, violating the fourth property in Definition 2.
@@ 3. Incremental Synthesis with Counting Constraints
The key advantage of counting constraints for synthesis is their monotonicity property: If $EGO$ has a strategy, such that $EGO$ plays an action $a$ at most $k$ times in $l$ turns (i.e. the strategy satisfies $CC_{max}(EGO,a,k,l)$), then $EGO$ also plays $a$ at most $k$ times in l1l-1 turns (i.e. the strategy satisfies $CC_{max}(EGO,a,k,l-1)$). In other words: The existence of a winning strategy for a game with counting constraint $CC_{max}(EGO,a,k,l-1)$ is a necessary condition for the winning strategy for a game with $CC_{max}(EGO,a,k,l)$. Moreover, only a strategy that fulfills $CC_{max}(EGO,a,k,l-1)$ can also fulfill $CC_{max}(EGO,a,k,l)$. From an algorithmic perspective, it is more favorable to search for strategies that satisfy $CC_{max}(EGO,a,k,l-1)$ then for strategies that satisfy $CC_{max}(EGO,a,k,l)$, since the graph that encodes the first (shorter) constraint is smaller than the one that encodes the latter (longer) constraint. Intuitively, this is caused by more memory that is needed for remembering the last $l$ own turns instead of only l1l-1 turns. The synthesis idea is related to the incremental approach used by synthesis with antichains [FJR09].
For a counting constraint of the form $CC_{min}(EGO,a,k,l-1)$ (EGOEGO plays action $a$ at least $k$ times out of l1l-1 of its turns”), we can conduct that if a strategy fulfills the constraint, it automatically also fulfills the larger constraint $CC_{min}(EGO,a,k,l)$. Hence, if we already have a strategy that fulfills $CC_{min}(EGO,a,k,l-1)$, it is needless to do the more challenging search for a strategy that fulfills $CC_{min}(EGO,a,k,l)$.
@@
Let $G=(A,\text{Win},CC)$ be a two-player game with counting constraints. The implications as shown below hold.
The player $EGO$ plays $a$ at most $k$ times out of $l$ of its own turns. The player $EGO$ plays $a$ at most lk+1l-k+1 times out of $l+1$ of its own turns. The player $EGO$ plays $\neg a$ at least $k$ times out of $l$ of its own turns. The player $EGO$ plays $\neg a$ at least $k$ times out of $l+1$ of its own turns.
Combining Theorem 1 and Lemma 2 yields for later increments: Instead of directly considering a counting constraint of the form “The player $EGO$ plays $a$ at most $k$ times out of $l$ of its own turns.”, we can start with “The player $EGO$ plays $\neg a$ at least lkl-k times out of lkl-k of its own turns.”, incrementally increase the constraint length to “The player $EGO$ plays $\neg a$ at least lkl-k times out of $l$ of its own turns.”, which translated back to the original counting constraint.
@@
Figure 4 shows a small part of the situation graph for the game introduced in Example 2 and Figure 1. A state from the game graph in Figure 1 combined with a history (presented as table in Figure 4) is a state “situation” of the situation graph. The left-most state of Figure 4 can be interpreted as follows: When this state is reached in the situation graph, a corresponding play in the original game is in state 3. In $EGO$s last turn, the player did not satisfy the action in counting constraint $C_{1}=$ “The player $EGO$ plays $x$ at least $2$ times out of $4$ of its own turns.” (“0” as first entry in the $C_{1}$-column), but satisfied the actions in $C_{2}$ and $C_{3}$ (“1” as first entry in the $C_{2}$- and $C_{3}$-columns). In the second to last turn of $EGO$, all of the actions in the $EGO$-constraints were fulfilled (“1” as second entry in the $C_{1}$-, $C_{2}$- and $C_{3}$-columns). For each counting constraint, only the amount of last turns relevant for the constraint are considered in the situation graph. Hence, some of the entries in the tables of Figure 4 are left empty. It can be seen that the leftmost and the uppermost situation fulfill all counting constraints. All other situations violate the constraint $C_{3}=$ “The player $EGO$ plays $y$ at most $2$ times out of $3$ of its own turns.”, since $EGO$ played $y$ in all of its last three turns (three times “1”-entries in the $C_{3}$-columns). Hence, no winning strategy of $EGO$ will suggest to play $x,y$ in state 3 of the game in Figure 1 when the state was reached as induced by the history in the leftmost situation of Figure 4.
@@
[Situation Graph] Let $G=(A,\text{Win},CC)$ a game with counting constraints with $A=(S,s_{0},S_{EGO},S_{ALTER},\Sigma_{EGO},\Sigma_{ALTER},\rightarrow)$. To ease the following notations, fix some order in the counting constraints $C=\{CC_{1},\dots,CC_{n}\}$.
For each counting constraint $CC_{m}(p,a,k,l)$ with $m\in\{min,max\}$, $p\in\{EGO,ALTER\}$, $a$ being a logical formula over $(\Sigma_{p})$ as defined in Definition 2 , $\in\mathbb{N}, klk\leq l$ define a mapping
@@
A situation is a tuple $(s,H)$ with $s\in S$ being a state in $A$, $H\in{}_{i=1}^{n}\text{codom}(h_{C_{i}})$ and $\text{codom}(f)=Y$ denoting the codomain of a function $f\colon X\rightarrow Y$. Denote the set of all situations by $\tilde{S}$, $\tilde{S}_{EGO}$ the set of situations with state (in $G$) controlled by $EGO$ and $\tilde{S}_{ALTER}$ the set of situations with a state controlled by $ALTER$. Define a mapping
@@
$|\{v^{\prime}_{j}\,|\,v^{\prime}_{j}\neq 0,j=1,\dots,l,\}|\quad\biggl\{\begin{aligned} \geq k,&\text{ if }m=min\\ \leq k,&\text{ if }m=max.\end{aligned}$
@@
The situation graph of a game is a deterministic Büchi automaton that represents the full specification of $EGO$, if the complete set of counting constraints is considered. Hence, finding a winning strategy in the situation graph directly translates to finding a winning strategy on the original game. Counting constraints are expressible as (long) LTL-formulae, consequently, using the full situation graph for synthesis is generally only doable in time double-exponential in the size of the LTL-specification [PR89]. The incremented synthesis approach avoids to construct the full situation graph. In fact, we will later use a slightly modified version of the situation graph that allows us to actually apply any synthesis algorithm that is also applicable on the original game $G$ without counting constraints.
[…]
  • The counting constraints in $\widetilde{CC}$ of the form $CC_{min}(EGO,a,k,l)$ are the same as the ones in $CC$ except the fact that they can be longer and the history-vector $\tilde{H}_{\pi(i)}$ is for such a constraint differs from $H_{i}$ only by potentially having additional entries at the end: $\forall i\in I:\,CC_{i}=CC_{min}(EGO,a_{i},k_{i},l_{i})\implies\widetilde{CC}_{\pi(i)}=CC_{min}(EGO,a_{i},k_{i},\tilde{l}_{i})$ with $l_{i}\leq\tilde{l}_{i}$ and $\tilde{H}_{\pi(i)}=(H_{i},X_{i})^{T}$ for some (arbitrary) $X_{i}\in\{0,1,none\}^{*}$
[…]
@@
In Figure 5, one of the situations of the previous example (Example 3, leftmost situation of Figure 4) is shown with one extension of the situation and one situation that it extends itself.
@@
For the scope of this article, we restrict ourselves to increment over constraints of the form $CC_{min}(EGO,a,k,l)$ and the usage of Lemma 2. The overall synthesis procedure is sketched in Figure 6. We explain the different phases of the incremental synthesis algorithm for a two-player finite-state game with counting constraints $G=(A,\text{Win},CC)$ based on this figure. We assume that the winning condition Win is either that of a safety, reachability, Büchi, Co-Büchi or parity game. The winning condition influences how the situation graph is modified in each increment. However, the approach might be applicable for more types of winning conditions.
@@
  • No outgoing transitions of situations that violate any constraint of $EGO$ in $\mathcal{C}$ are considered. If a situation violates any constraint of player $EGO$, the situation will not be in the winning region of the situation graph. Hence, for constructing a winning strategy for $EGO$ in the situation graph (which directly translates to a potentially non-positional winning strategy of $EGO$ in $G$), outgoing transitions of a situation with violated $EGO$-constraints are not relevant for finding winning strategies in the current increment. Note that an extension of a situation that violate a constraint can again satisfy a constraint. Therefore, such extensions are considered again in later increments.
  • Situations that violate any constraint of $ALTER$ are not added to the situation graph. For any situation that satisfies all constraints of player $A$ and is controlled by $ALTER$, ALTERALTER can choose an action as next turn, such that the reached situation also satisfies all of its counting constraints (as a consequence of the last property in Definition 2 and Theorem 1). Hence, all relevant behavior of $ALTER$ is still captured when omitting situations that violate $ALTER$-constraints.
@@
The third criterion on not extending the situation graph from extensions of situations of the winning region of a previous increment is a key factor in reducing the required computational time and memory for incremental synthesis. While the increasing lengths of the considered counting constraints extends the lengths of the relevant histories and the number of potentially reachable situations, avoiding to construct parts of the situation graph by using knowledge retrieved in previous increments prunes the graph again. Figures 7(a) and 7(b) illustrate the difference between the definition of the situation graph (Definition 4) and the modified version in this step.
@@
[…]
Parity winning condition $\text{Win}=\mathtt{Parity}(\Omega)$: For parity winning conditions, we start by defining the coloring of situations in $Sit_{\mathcal{C}}$. Denote by $\Omega\colon S\to\mathbb{N}_{0}$ the coloring function in $(A,\text{Win})$. Define
[…]
as coloring on $Sit_{\mathcal{C}}$. Situations (except for the new sink-locations) inherit their color from their state in $S$. Whenever a play enters the winning sink $W$, EGOEGO will win the play, since each situation in $W$ is visited infinitely often and is colored with 0, being the minimal coloring and even. Similarly, if a play enters the sink $L$, EGOEGO is going to loose, since $L$ will never be left, hence only finitely many visited situations are not in $L$ and all situations in $L$ are colored with the uneven color 1.
Reachability winning condition $\text{Win}=\mathtt{Reach}(R)$: Compared to the other presented winning conditions, adapting the winning condition and the situation graph for being solvable by an algorithm for reachability games is more complex. This is due to the fact that all of the other considered winning conditions share the “infinity aspect” with the condition of satisfying a set of window counting constraints, but for winning a play of a reachability game, it suffices to visit one of the states in $R$ just once. For a reachability game, the construction of $\widetilde{\text{Win}}$ as for the other winning conditions would lead to plays that visit any situation $(s,H)\in\tilde{S}$ with $s\in R$ and violating a counting constraint of $EGO$ later as being in the winning region. To circumvent this issue, it is possible to play a safety game first and play the reachability game only on the winning region of the safety game:
@@
  • Alternating: In this mode, the constraints are extended in turns. Again, choose a fixed order $C_{1},\dots,C_{m}$ of the counting constraints of the form $CC_{min}(EGO,a,k,l)$ in $\mathcal{C}$. In the first increment, the first constraint that is not yet completely extended is extended by one. For following increments: If in the last increment constraint $C_{i}$ was extended, extend now the next constraint after $C_{i}$ that has not yet reached its full length as in $\mathscr{C}$ (consider the ordering as circular, i.e.modulo $m$).

### v3

@@ 2. Games with Counting Constraints
    $CC_{min}(p,a,k,l)$ is defined as the abbreviation for “The player $p$ plays $a$ at least $k$ times out of $l$ of its own turns.”. $CC_{min}(p,a,k,l)$ is satisfied on $\pi$, if for all $i\in\mathbb{N}_{0}$ holds $|\{m\in\mathbb{N}_{0}\,|\,u^{p}_{m}\models a,\,i\leq m\leq i+l\}|\geq k$.
@@
  • “The player $EGO$ plays $y$ at least $2$ times out of $5$ of its own turns.” and
@@
The play $\pi=(1,\{y\},2,\{a\},3,\{x\},2,\{a\},3,\{x,y\},4,\{a\})^{\omega}$ on the game graph satisfies all of the four window counting constraints, since each “sliding window” shows constraint satisfaction. Figure 2 illustrates the constraint fulfillment for the first two constraints.
The prefix $pre=(1,\{x\},4,\{a\},1,\{x\},4,\{x\},1,\{x\},4)$ of another play on the shown graph is not satisfying “The player $EGO$ plays $y$ at least $2$ times out of $5$ of its own turns.”, since no matter what $EGO$ does in its next move (even if it would be allowed to change the game graph for the next turns), it is not possible to extend prepre to a play that satisfies the constraint.
@@
  • Win is a winning condition for $EGO$ on the game graph $A$.
  • CCCC is a finite set of counting constraints.
@@
Player $EGO$ wins a play on $G$, if $ALTER$ violates one of its constraints in the play or if Win is fulfilled and $EGO$ satisfies all of its counting constraints on the play. A strategy $\varphi$ of $EGO$ is winning for $EGO$ (or a winning strategy of $EGO$), if $EGO$ wins all plays in $O(G,\varphi)$.
The fourth property of a game with counting constraints seems counter-intuitive, since it restricts the considered games to those in which $ALTER$ can basically ensure in each step to satisfy all of its constraints without any need of planning ahead. The main reason for introducing this property is as follows: We are not interested in plays that are only won by $EGO$ due to forcing $ALTER$ to lose due to violated counting constraints. Dismissing the property would lead to situations like depicted in Figure 3: The $EGO$-player does not have any counting constraints to obey but needs to avoid an unsafe state (colored in gray). $ALTER$ has the counting constraint ALTERALTER plays $b$ at least 1 time in 1 turn.” (hence, $ALTER$ needs to play $b$ in each of its turns). In the initial state, choosing to enter the upper sub-graph lets $EGO$ win. However, taking the transitions to the lower subgraph and consequently visiting an unsafe state would also lead to a win for $EGO$, since $ALTER$ is violating its constraint in the resulting play. This problem of winning by falsifying the assumptions, but not leading to intended behavior of the player is well-known [15], [1]. The fourth property of Definition 2 is our way to exclude such trivial solutions inherent to the problem space of synthesis. The setting fits especially for applications in which the behavior of $EGO$ is influenced by the behavior of the environment (represented by $ALTER$), but not the other way around. An example for that is a transport robot on a factory floor that is supposed to transport materials to and from machines. Those machines usually have a buffer for materials, such that their behavior is independent of the robot (assuming that an empty input buffer or overfull output buffer would be preceded by a violation of the robots specification). However, which transport tasks need to be done by the robot highly depends on the machines.
@@
Figure 3. Graph for a safety game that cannot be used in a game with the counting constraint ALTERALTER plays $b$ at least 1 time in 1 turn.”. Circles represent states controlled by $EGO$, diamond-shaped states are controlled by $ALTER$. The state colored in gray marks an unsafe state. When the lower subgraph is entered, ALTERALTER cannot fulfill its counting constraint, violating the fourth property in Definition 2.
@@
The following lemma characterizes the size of a game graph encoding the combination of a (smaller) graph and a set of counting constraints.
Lemma 1.
Let $G=(A,\text{Win},CC)$ be a game with counting constraints. Let $A^{\prime}$ be the expanded game graph that encodes the combination of the game graph $A$ with the set $CC$ of counting constraints. The number of states of the game graph $A$ is exponential in the sizes of the counting constraints in $CC$, where the size of a constraint $CC_{m}(p,a,k,l)$, $m\in\{min,max\}$, is the magnitude of the parameter $l$.
Proof 2.1.
For a constraint of the form $CC_{min}(p,a,k,l)\in CC$, each state $s$ in $A$ that is controlled by player $p$ is enriched with information on when in the last $l$ turns the action $a$ was played. Assume this information is encoded as binary sequence of length $l$, in which a “1” in the $i$th position expressing that action $a$ was played $i$ turns ago ($1\leq i\leq l$). There exist $\sum_{n=k}^{l}\binom{l}{n}$ different sequences of the length $l$ that represent the constraint fulfillment for the last $l$ turns. Thus each state $s$ of the original game graph $A$ is refined into $\sum_{n=k}^{l}\binom{l}{n}$ states in the combined game graph $A^{\prime}$ (plus some more states for encoding the beginning of the game, in which there are not yet $l$ turns played). For a counting constraint of the form $CC_{max}(p,a,k,l)$, a state in $A$ is refined into $\sum_{n=0}^{k}\binom{l}{n}$ states in $A^{\prime}$. For a set of constraints, the number of states in $A^{\prime}$ that refine a state in $A$ is the product of the individual sums of each constraint and the number of states in the original graph. Hence, the number of states in $A^{\prime}$ is roughly exponential in the size of the product of all constraint lengths, leading to exponential costs for strategy synthesis on $A^{\prime}$.
Note that for some games, a memory-efficient winning strategy for the player $EGO$ exists. For example, a winning strategy for a game with trivial game graph graph $A$ in which $EGO$ can play one of its actions in each turn without further restriction combined with the constraints $CC_{min}(EGO,a,k,l)$ and $CC_{min}(EGO,b,l-k,l)$ can be realized by first playing $a$ in $k$ turns, followed by playing $b$ in the next lkl-k turns, requiring only one counter over $l$ values. The combination of counting constraints with a non-trivial game graph raises the need for more complex strategies. Forward-synthesis algorithms aim to avoid the full expansion of the graph, heuristically leading to savings in the synthesis costs. The following section explains how the structure of counting constraints can contribute to this goal.
@@ 3. Incremental Synthesis with Counting Constraints
The key advantage of counting constraints for synthesis is their monotonicity property: If $EGO$ has a strategy, such that $EGO$ plays an action $a$ at most $k$ times in $l$ turns (i.e. the strategy satisfies $CC_{max}(EGO,a,k,l)$), then $EGO$ also plays $a$ at most $k$ times in l1l-1 turns (i.e. the strategy satisfies $CC_{max}(EGO,a,k,l-1)$). In other words: The existence of a winning strategy for a game with counting constraint $CC_{max}(EGO,a,k,l-1)$ is a necessary condition for the winning strategy for a game with $CC_{max}(EGO,a,k,l)$. Moreover, only a strategy that fulfills $CC_{max}(EGO,a,k,l-1)$ can also fulfill $CC_{max}(EGO,a,k,l)$. From an algorithmic perspective, it is more favorable to search for strategies that satisfy $CC_{max}(EGO,a,k,l-1)$ then for strategies that satisfy $CC_{max}(EGO,a,k,l)$, since the graph that encodes the first (shorter) constraint is smaller than the one that encodes the latter (longer) constraint. Intuitively, this is caused by more memory that is needed for remembering the last $l$ own turns instead of only l1l-1 turns. The synthesis idea is related to the incremental approach used by synthesis with antichains [10].
For a counting constraint of the form $CC_{min}(EGO,a,k,l-1)$ (EGOEGO plays action $a$ at least $k$ times out of l1l-1 of its turns”), we can conclude that if a strategy fulfills the constraint, it automatically also fulfills the larger constraint $CC_{min}(EGO,a,k,l)$. Hence, if we already have a strategy that fulfills $CC_{min}(EGO,a,k,l-1)$, there is no need to search for a strategy that fulfills $CC_{min}(EGO,a,k,l)$, which is more computationally demanding.
@@
Let $G=(A,\text{Win},CC)$ be a two-player game with counting constraints. For each (infinite) play on $G$ hold the following implications:
The player $EGO$ plays $a$ at most $k$ times out of $l$ of its own turns. The player $EGO$ plays $a$ at most $k+1$ times out of $l+1$ of its own turns. The player $EGO$ plays $\neg a$ at least lkl-k times out of $l$ of its own turns. The player $EGO$ plays $\neg a$ at least lkl-k times out of $l+1$ of its own turns.
Combining Theorem 2 and Lemma 3 yields for later increments: Instead of directly considering a counting constraint of the form “The player $EGO$ plays $a$ at most $k$ times out of $l$ of its own turns.”, we can start with “The player $EGO$ plays $\neg a$ at least lkl-k times out of lkl-k of its own turns.”, incrementally increase the constraint length to “The player $EGO$ plays $\neg a$ at least lkl-k times out of $l$ of its own turns.”, which translates back to the original counting constraint.
@@
Figure 4 shows a small part of the situation graph for the game introduced in Example 2 and Figure 1. A state from the game graph in Figure 1 combined with a history (presented as table in Figure 4) is a situation of the situation graph. The left-most state of Figure 4 can be interpreted as follows: When this state is reached in the situation graph, a corresponding play in the original game is in state 3. In $EGO$s last turn, the player did not satisfy the action in counting constraint $C_{1}=$ “The player $EGO$ plays $x$ at least $2$ times out of $4$ of its own turns.” (“0” as first entry in the $C_{1}$-column), but satisfied the actions in $C_{2}$ and $C_{3}$ (“1” as first entry in the $C_{2}$- and $C_{3}$-columns). In the second to last turn of $EGO$, all of the actions in the $EGO$-constraints were fulfilled (“1” as second entry in the $C_{1}$-, $C_{2}$- and $C_{3}$-columns). For each counting constraint, only the amount of last turns relevant for the constraint are considered in the situation graph. Hence, some of the entries in the tables of Figure 4 are left empty. It can be seen that the leftmost and the uppermost situation fulfill all counting constraints. All other situations violate the constraint $C_{3}=$ “The player $EGO$ plays $y$ at most $2$ times out of $3$ of its own turns.”, since $EGO$ played $y$ in all of its last three turns (three times “1”-entries in the $C_{3}$-columns). Hence, no winning strategy of $EGO$ will suggest to play $x,y$ in state 3 of the game in Figure 1 when the state was reached as induced by the history in the leftmost situation of Figure 4.
@@
[Situation Graph] Let $G=(A,\text{Win},CC)$ a game with counting constraints with $A=(S,s_{0},S_{EGO},S_{ALTER},\Sigma_{EGO},\Sigma_{ALTER},\rightarrow)$. To ease the following notations, fix some order in the counting constraints $CC=\{CC_{1},\dots,CC_{n}\}$.
For each counting constraint $C=CC_{m}(p,a,k,l)$ with $m\in\{min,max\}$, $p\in\{EGO,ALTER\}$, $a$ being a logical formula over $\Sigma_{p}$ as defined in Definition 2, $\in\mathbb{N}, klk\leq l$ define a mapping
@@
A situation is a tuple $(s,H)$ with $s\in S$ being a state in $A$, $H\in\prod_{i=1}^{n}\text{codom}(h_{C_{i}})$ and $\text{codom}(f)=Y$ denoting the codomain of a function $f\colon X\rightarrow Y$. Denote the set of all situations by $\tilde{S}$, $\tilde{S}_{EGO}$ the set of situations with state (in $G$) controlled by $EGO$ and $\tilde{S}_{ALTER}$ the set of situations with a state controlled by $ALTER$. Define a mapping
@@
$|\{j\in\{1,\dots,l\}\,|\,v^{\prime}_{j}\neq 0\}|\quad\biggl\{\begin{aligned} \geq k,&\text{ if }m=min\\ \leq k,&\text{ if }m=max.\end{aligned}$
@@
The incremental synthesis approach avoids to construct the full situation graph, which is exponential in the constraint lengths (Lemma 1, as the expanded game graph in the lemma is basically a situation graph without details on the notation). In fact, we will later use a slightly modified version of the situation graph that allows us to actually apply any synthesis algorithm that is also applicable on the original game $G$ without counting constraints.
The general idea is to start with a rather small game by using counting constraints of small lengths and increment over the length. In each increment, a part of the corresponding situation graph is constructed and analyzed and knowledge that can be reused in following increments is identified. This knowledge is determining which parts of the situation graph for the next increment needs to be constructed and which parts can be omitted, relying on 2. To this end, we introduce extensions of situations in order to compare situations from one increment with situations in another increment. {defi}[Extensions of situations] Let $(s,H)$ be a situation with counting constraints $CC$, $H=(H_{1},\dots,H_{n})$, $H_{i}$ representing the history for counting constraint $CC_{i}\in CC$.
  • A set $CC^{\prime}$ of counting constraints is an extension of counting constraints $CC$, if the only difference between $CC$ and $CC^{\prime}$ are longer counting constraints of the form $CC_{min}(EGO,a,k,l)$. Formally, $CC^{\prime}$ is an extension of $CC$, if $|CC|=|CC^{\prime}|$ and if for all $CC_{i}\in CC$, there exists $CC_{i}^{\prime}\in CC^{\prime}$ with
    $CC_{i}^{\prime}=\biggl\{\begin{aligned} CC_{min}(EGO,a,k,l^{\prime})\text{ with }l^{\prime}\geq l,&\text{ if }CC_{i}=CC_{min}(EGO,a,k,l)\\ CC_{i},&\text{ else.}\end{aligned}$
  • A history $H^{\prime}$ with counting constraints $CC^{\prime}$ is an extension of a history $H$, if $CC^{\prime}$ is an extension of $CC$ and $H^{\prime}$ differs from $H$ only by additional entries for constraints of the form $CC_{i}=CC_{min}(EGO,a,k,l)$. Formally: $H^{\prime}$ is an extension of $H$, if the corresponding counting constraints $CC^{\prime}$ are an extension of $CC$ and it exists an $X\in(\{0,1,\text{none}\}^{*})^{|CC|}$, such that $H^{\prime}=H.X$ (element-wise concatenation).
  • A situation $(s^{\prime},H^{\prime})$ is an extension of a situation $(s,H)$, if $s=s^{\prime}$ and $H^{\prime}$ is an extension of $H$. In a nutshell, an extension $(s^{\prime},H^{\prime})$ of a situation $(s,H)$ provides more information than $(s,H)$ on former turns of $EGO$ with respect to constraints of the form $CC_{min}(EGO,a,k,l)$ and equals $(s,H)$ in all other aspects.
@@
In Figure 5, one of the situations of the previous example (Example 3, leftmost situation of Figure 4) is shown with one extension of the situation and one situation that extends itself.
@@
For the scope of this article, we restrict ourselves to increment over constraints of the form $CC_{min}(EGO,a,k,l)$ and the usage of Lemma 3. The overall synthesis procedure is sketched in Figure 6. We explain the different phases of the incremental synthesis algorithm for a two-player finite-state game with counting constraints $G=(A,\text{Win},CC)$ based on this figure. In each increment, a synthesis algorithm of the readers choice that can determine the winning region of a game $(A,\text{Win})$, with the chosen winning condition is used. Since the concept of a winning region only makes sense for positionally determined games, we assume that the winning condition Win is either that of a safety, reachability, Büchi, Co-Büchi or parity game. The winning condition influences how the situation graph is modified in each increment. We checked the modification explicitly for the mentioned winning conditions. However, the approach might be applicable for more types of winning conditions, especially for other positionally determined games.
@@
  • No outgoing transitions of situations that violate any constraint of $EGO$ in $\mathcal{C}$ are considered, since they are clearly not part of any winning strategy for $EGO$ in the game $(A,\text{Win},\mathcal{C})$. Note that an extension of a situation that violate a constraint can again satisfy a constraint. Therefore, such extensions are considered again in later increments.
  • Situations that violate any constraint of ALTERALTER are not added to the situation graph. This is not leading to any relevant restriction on the behavior of ALTERALTER due to the last property in Definition 2 and Theorem 1.
The first criterion on not extending the situation graph from extensions of situations of the winning region of a previous increment is a key factor in reducing the required computational time and memory for incremental synthesis. While the increasing lengths of the considered counting constraints extends the lengths of the relevant histories and the number of potentially reachable situations, avoiding to construct parts of the situation graph by using knowledge retrieved in previous increments prunes the graph again. Figures 7(a) and 7(b) illustrate the difference between the definition of the situation graph (Definition 4) and the modified version in this step.
@@
Reachability winning condition $\text{Win}=\mathtt{Reach}(R)$: Compared to other winning conditions like safety, Büchi, co-Büchi or parity, adapting the winning condition and the situation graph for being solvable by an algorithm for reachability games is more complex. This is due to the fact that all of the other considered winning conditions share the “infinity aspect” with the condition of satisfying a set of window counting constraints, but for winning a play of a reachability game, it suffices to visit one of the winning states in $R$ just once. For a reachability game, the construction of $\widetilde{\text{Win}}$ as for the other winning conditions would lead to plays that visit any situation $(s,H)\in\tilde{S}$ with $s\in R$ and violating a counting constraint of $EGO$ later as being in the winning region. To circumvent this issue, it is possible to play a safety game first and play the reachability game only on the winning region of the safety game:
@@
  • Round-robin: In this mode, the constraints are extended in turns. Again, choose a fixed order $C_{1},\dots,C_{m}$ of the counting constraints of the form $CC_{min}(EGO,a,k,l)$ in $\mathcal{C}$. In the first increment, the first constraint that is not yet completely extended is extended by one. For following increments: If in the last increment constraint $C_{i}$ was extended, extend now the next constraint after $C_{i}$ that has not yet reached its full length as in $\mathscr{C}$ (consider the ordering as circular, i.e.modulo $m$).